OpenStudy (anonymous):

Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x(to the 2nd power) − 1. Calculate f(3) and g(3), then multiply them together. Calculate f(3) and g(3), then write the results as an ordered pair. Calculate f(3) and g(3), then subtract them. Calculate g(3), plug the result into f(x) and simplify.

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

\[f(3) = 3(3) - 2 = 9 - 2 = 7; g(3) = 3^2 = 9; (g(f(x)) = 63\] Thats for the first.